OpenStudy (australopithecus):

Determine whether the series converges or diverges. Verify that the test used is applicable: I have a) \[\sum_{n=2}^{\infty} \frac{121+\sin(121n)}{n^{21} + 21n}\] b) \[\sum_{n=4}^{\infty}\frac{2012}{\ln(n^{n})}\] a) I split the series and used the comparison test \[\frac{121}{n^{21}+21n}\le \frac{121}{n^{21}}\] \[\frac{\sin(121n)}{n^{21}+21n} \le \frac{1}{n^{21}}\] by p test they both converge b) I used the integral test, determined -the series is decreasing \[\frac{2012}{\ln(n^{n})} \ge \frac{2012}{\ln(n+1)^{n+1}}\] -the function is decreasing \[f(x) = -\frac{2012}{\ln(n^{

OpenStudy (australopithecus):

-the function is decreasing -2012/ln(n^(n))^(2) - and the function is always positive thus I applied the integral test \[\int\limits_{\infty}^{4}\frac{2012}{\ln(n^{n})} = 2012\ln|n^{n}|\left(\begin{matrix}\infty \\ 4\end{matrix}\right)\]

OpenStudy (australopithecus):

please ignore the brackets I don't know how to use proper integral notation on this site, anyways, I got that the integral was equal to infinity and thus divergent therefore the series is divergent by integral test. My main question is did I do this correctly and how do I verify that the test used is applicable?

OpenStudy (australopithecus):

wait I made a mistake solving the integral of 121/ln(x^(x))

OpenStudy (australopithecus):

u = ln(x) + 1 du/(ln(x)+1) = dx

OpenStudy (australopithecus):

this integral is not easily solved ugh

OpenStudy (australopithecus):

nvm I will just use partial fractions :)